OpenStudy (anonymous):

Find three consecutive positive integers such that the square of the first, increased by the last, is 22

OpenStudy (anonymous):

If 3x3 is 9, and 9 plus 13=22, then the answer is 3, 6, and 9

OpenStudy (anonymous):

thx

OpenStudy (radar):

Let x = first positive integer Let x + 1 = second positive integer Let x + 2 = third positive integer. Then as problem states \[x ^{2} + (x + 2)=22\] \[x ^{2}+x-20 =0\](after simplifing. factor quadratic getting (x+5)(x-4) = 0 Select positive value of x or x-4=0, thus x=4 (first positive integer x + 2=4 + 2 = 6 is the third positive integer: 4, 5, and 6

OpenStudy (anonymous):

thank you

OpenStudy (radar):

To verify: square the first positive integer: 4 squared (4X 4) is 16 Third integer is 6 16 + 6 = 22

OpenStudy (anonymous):

thanks a million